# EGG-SR: Embedding Symbolic Equivalence into Symbolic Regression via Equality Graph

### v1

@@ A.3 Implementation of E-Graph Functionalities
def apply_rewrite_rules(eg: EGraph, rules: List[Rule]):
eclasses = eg.eclasses()
matches = []
for rule in rules:
for eid, env in ematch(eclasses, rule.lhs):
matches.append((rule, eid, env))
for rule, eid, env in matches:
new_eid = eg.substitute(rule.rhs, env)
eg.merge(eid, new_eid)
eg.rebuild()
[…]
eg = EGraph(rules_to_s_expr(fn))
for it in range(max_iter):
apply_rewrite_rules(eg, rules)
We employ the Graphviz API to visualize the e-classes, e-nodes, and their connections. Several visualization example is shown in section F.
A.4 Expression Extraction
@@
The algorithm iteratively updates costs for all e-nodes until convergence, i.e., when no further changes occur. This method correctly handles shared subexpressions and potential cycles within the e-graph.
(1) Cost-based extraction, which retrieves a simplified expression by minimizing a user-defined cost function over operators;
(2) Random walk-based sampling, which draws a batch of expressions randomly from the e-graph. Starting from an e-class with no incoming edges, we randomly select an e-node within the current e-class and transition to the e-classes connected to that e-node. This process continues until an e-node with no outgoing edges is reached. The obtained sequence of visited e-nodes corresponds to drawing a valid expression from the constructed e-graph.
A.5 Implementation Remark
@@
Appendix B Proof of Theoroem 3.1
B.1 Problem settings
Necessary Definitions of Markov Decision Process Consider a deterministic, finite-horizon Markov Decision Process (MDP) with state space $\mathcal{S}$ and action space $\mathcal{A}$. At each stage $t\in\{0,1,\dots,H-1\}$ the agent observes its current state $s_{t}\in\mathcal{S}$, selects an action $a_{t}\in\mathcal{A}$, and deterministically transitions to the next state $s_{t+1}=f(s_{t},a_{t})$ while receiving a bounded reward $r_{t}\in[0,1]$. Let $H$ denote the maximum planning horizon. The total discounted reward over the $H$-step horizon is $\sum_{t=0}^{H-1}\gamma^{t}r_{t}$, where $\gamma\in(0,1)$ is the discount factor. For any state–action pair $(s,a)$, the state–action value is defined as $Q(s,a)\coloneqq\max_{\pi}\mathbb{E}_{\tau\sim\pi}\!\left[\sum_{t=0}^{H-1}\gamma^{t}r_{t}\middle|s_{0}=s,\,a_{0}=a\right]$, where $\pi$ is any policy and $\tau=(s_{0},a_{0},\dots,s_{H-1},a_{H-1})$ is a trajectory generated by following $\pi$. Because the MDP is deterministic, the expectation is taken only over the agent’s policy-induced action sequence. The optimal value of a state $s$ is $V(s)\coloneqq\max_{a\in\mathcal{A}}Q(s,a)$.
Connection to Symbolic Regression. In our symbolic regression setting, a state corresponds to a sequence of production rules that define a partially constructed expression, or equivalently, to the class of mathematical expressions that can be generated by completing this partial structure. An action represents the selection of an available production rule that extends the current expression. The state space is the set of all possible expressions of maximum length $H$, and the action space is the set of production rules.
A planning algorithm is any procedure that, given a model of the MDP and a limited computational budget, uses simulated rollouts to estimate the value of available actions and to recommend a single action for execution. Suppose that after $n$ simulations of a planning algorithm, the agent recommends an action $a_{T}$ to execute at the current state $s$. Its estimated value is $Q(s,a_{T})$, obtained by evaluating the discounted return that results from first taking $a_{T}$ and then following an optimal policy for the remaining steps. The simple regret of this recommendation is
$\mathtt{regret}(T)\coloneqq V(s)-Q(s,a_{T}),$ (7)
which quantifies the loss in discounted return incurred by choosing $a_{T}$ instead of the truly optimal action. A smaller $\mathtt{regret}(n)$ indicates that the planning algorithm used its limited simulation Budget more effectively to identify a near-optimal action.
@@ B.1 Problem settings
In standard MCTS, the root of the search tree corresponds to the initial state $s_{0}\in S$. A node at depth $h$ represents an action sequence $a=(a_{1},\ldots,a_{h})\in\mathcal{A}^{h}$, which leads to a terminal state $s(a)$. The search tree at iteration $n$ is denoted $\mathtt{tree}(T)$, and its maximum expanded depth is $d_{n}$.
Historical research has shown that the UCT principle suffers from a theoretical worst case in which it is difficult to derive meaningful regret bounds, making it unsuitable for analyzing the benefit of Egg over standard MCTS. Therefore, we assume that MCTS operates under the Optimistic Planning for Deterministic Systems (OPD) framework (DBLP:conf/acml/LeurentM20) rather than the UCT principle.
@@
$\mathtt{regret}(T)=\widetilde{\mathcal{O}}\!\left(n^{-\frac{\log(1/\gamma)}{\log\kappa}}\right).$ (9)
Here, the soft-O version of big-O notation is defined as: $f_{n}=\widetilde{\mathcal{O}}(n^{-\alpha})$ means that decays at least as fast as $n^{-\alpha}$, up to a polylog factor.
@@
The theoretical analysis of Egg-MCTS is built on top of the analysis procedure in (DBLP:conf/acml/LeurentM20). Unlike their analysis, which requires unrolling the graph into a tree and addressing potentially infinite-length paths induced by cycles, our analysis starts from the unrolled tree derived from the graph in (DBLP:conf/acml/LeurentM20).
@@
Let $\mathtt{tree}(T)$ be the search tree after $T$ expansions. Each node corresponds to an action sequence $(a_{0},a_{1},\dots,a_{H-1})$ from the root, and let $s$ denote the MDP state reached by executing this sequence from the root state $s_{0}$. The true value associated with this node is the optimal value of the reached state, denoted by $V(s)$. A pair of functions $(L_{T},U_{T})$ is said to provide lower and upper bounds for $V$ on $\mathtt{tree}(T)$ if
[…]
@@
We define the near-optimal branching factor associated with bounds $(L_{T},U_{T})$ as
$\kappa(L_{T},U_{T})\coloneqq\limsup_{h\to\infty}\big|T_{h}^{\infty}(L_{T},U_{T})\big|^{1/h}\in[1,K],$ (10)
@@
Since $(L_{T},U_{T})$ tighten with $T$, the sequence of bounds $\bigl(L_{T},U_{T}\bigr)_{T\geq 0}$ is non-increasing, i.e.,
$0\leq\cdots\leq L_{T-1}\leq L_{T}\leq V\leq U_{T}\leq U_{T-1}\leq\cdots\leq V_{\max}$
they will finally converges to a limit $\kappa_{\infty}=\lim_{n\to\infty}\kappa(L_{T},U_{T})$.
@@
Let $\kappa_{T}=\kappa(L_{T},U_{T})$, and define $\kappa_{\infty}=\lim_{n\to\infty}\kappa(L_{T},U_{T})\in[1,K]$. Then Egg-MCTS achieves the regret bound
@@
The analysis is similar to that of DBLP:conf/acml/LeurentM20. In their approach, the graph is unrolled into a tree to leverage the analysis of Theorem B.1. This unrolled tree contains every action sequence that can be traversed in $G_{n}$. There would exist a path of infinite length if the graph contains cycles. The behavior of the graph $G_{n}$ is therefore analyzed through its corresponding unrolled tree $\mathtt{UnrollTree}(T)$.
Our Egg-MCTS behaves almost identically to this unrolled tree, except that it contains no paths of infinite length, since our algorithm cannot empirically generate or update along an infinite action sequence. Transporting the graph-based value bounds $(L_{T},U_{T})$ to the unrolled tree and applying the depth–regret argument of DBLP:conf/acml/LeurentM20 yields the rate in equation 11. The detailed derivation is therefore omitted here for brevity. ∎
@@
Appendix C Proof of Theorem 3.2
C.1 Notation and Assumptions
@@
The reward of the expression $\phi$ is defined as the goodness-of-fit on the dataset $D$.
C.2 Proof of Unbiasedness
@@
$\mathbb{E}\;\left[\,\nabla_{\theta}\log p_{\theta}(\tau)\,\middle|\,\phi\right]=\sum_{\tau\in\mathcal{S}_{\phi}}\nabla_{\theta}\log p_{\theta}(\tau)\;\frac{p_{\theta}(\tau)}{q_{\theta}(\phi)}=\frac{1}{q_{\theta}(\phi)}\sum_{\tau\in\mathcal{S}_{\phi}}\nabla_{\theta}p_{\theta}(\tau)$ $=\frac{1}{q_{\theta}(\phi)}\nabla_{\theta}\sum_{\tau\in\mathcal{S}_{\phi}}p_{\theta}(\tau)$
@@
where we used class-invariance of the reward and equation 12. This equals the expression obtained for $g_{\mathtt{egg}}$, proving unbiasedness. ∎
C.3 Proof of Variance Reduction
@@
$\mathbb{E}\;\left[\,Z\mid\phi\right]=\mathtt{reward}(\phi)\,\mathbb{E}\;\left[\,\nabla_{\theta}\log p_{\theta}(\tau)\mid\phi\right]=\mathtt{reward}(\phi)\,\nabla_{\theta}\log q_{\theta}(\phi)=g_{\mathtt{egg}}(\theta).$
Thus, $g_{\mathtt{egg}}(\theta)$ is the Rao–Blackwellization of $g(\theta)$ with respect to $\sigma(\phi)$ (8a19a051-2be1-32f8-8b83-075ff1f85473).
@@
$\mathrm{Var}\left(g_{\mathtt{egg}}(\theta)\right)\;\leq\;\mathrm{Var}\left(g(\theta)\right)$.
@@
$\mathrm{Var}(Z)=\mathrm{Var}\big(\mathbb{E}[Z\mid\phi]\big)\;+\;\mathbb{E}\big[\mathrm{Var}(Z\mid\phi)\big]\;\geq\;\mathrm{Var}\;\big(\mathbb{E}[Z\mid\phi]\big).$
Substituting $Z=g(\theta)$ and $\mathbb{E}[Z\mid\phi]=g_{\mathtt{egg}}(\theta)$ yields
$\mathrm{Var}\left(g(\theta)\right)\geq\mathrm{Var}\left(g_{\mathtt{egg}}(\theta)\right).$
@@
Appendix D Implementation of Egg-embedded Symbolic Regression
D.1 Implementation of Egg-MCTS
The algorithmic procedure of Egg-MCTS is described in Section 3.2. The code implementation is adapted from existing MCTS frameworks  [^4] https://github.com/isds-neu/SymbolicPhysicsLearner.

### v2

@@ A.3.1 Construction
1  def apply_rewrite_rules(eg: EGraph, rules: List[Rule]):2   eclasses = eg.eclasses()3   matches = []4   for rule in rules:5   for eid, env in ematch(eclasses, rule.lhs):6   matches.append((rule, eid, env))7   for rule, eid, env in matches:8   new_eid = eg.substitute(rule.rhs, env)9   eg.merge(eid, new_eid)10   eg.rebuild()11  12  def equality_saturation(fn, rules, max_iter=20):13   eg = EGraph(rules_to_s_expr(fn))14   for it in range(max_iter):15   apply_rewrite_rules(eg, rules)We employ the Graphviz API to visualize the e-classes, e-nodes, and their connections. Several visualization example is shown in section D.
A.3.2 Expression Extraction
@@
The algorithm iteratively updates costs for all e-nodes until convergence, i.e., when no further changes occur. This strategy yields short, simplified expressions and has been used in prior work [de2025equality, rEGGression], which employed e-graphs to simplify overly complex expressions in genetic programming. It is consistent with the philosophical principle of Occam’s razor, which favors simpler expressions over more complex ones.
Random walk-based sampling, which draws a batch of expressions randomly from the e-graph. Starting from an e-class with no incoming edges, we randomly select an e-node within the current e-class and transition to the e-classes connected to that e-node. This process continues until an e-node with no outgoing edges is reached. The obtained sequence of visited e-nodes corresponds to drawing a valid expression from the constructed e-graph.
A.3.3 Implementation Remark
Existing e-graph implementations are available in Rust (egg library, [DBLP:journals/pacmpl/WillseyNWFTP21]), Haskell (Hegg or srtree library), and Julia (Metatheory.jl), or wrappers for Python (egglog-python, [shanabrook2024egglogpythonpythoniclibrary, rEGGression]). While these frameworks offer rich APIs for diverse use cases, none is directly tailored to grammar-based symbolic expressions.
@@ A.3.3 Implementation Remark
Appendix B Theoretical Justification
B.1 Necessary Definitions of Markov Decision Process.
Consider a deterministic, finite-horizon Markov Decision Process (MDP) with state space $\mathcal{S}$ and action space $\mathcal{A}$. At each stage $t\in\{0,1,\dots,H-1\}$ the agent observes its current state $s_{t}\in\mathcal{S}$, selects an action $a_{t}\in\mathcal{A}$, and deterministically transitions to the next state $s_{t+1}=f(s_{t},a_{t})$ while receiving a bounded reward $r_{t}\in[0,1]$. Let $H$ denote the maximum planning horizon. The total discounted reward over the $H$-step horizon is $\sum_{t=0}^{H-1}\gamma^{t}r_{t}$, where $\gamma\in(0,1)$ is the discount factor. For any state–action pair $(s,a)$, the state–action value is defined as:
$Q(s,a)\coloneqq\max_{\pi}\mathbb{E}_{\tau\sim\pi}\!\left[\sum_{t=0}^{H-1}\gamma^{t}r_{t}\middle|s_{0}=s,\,a_{0}=a\right]$
where $\pi$ is any policy and $\tau=(s_{0},a_{0},\dots,s_{H-1},a_{H-1})$ is a trajectory generated by following $\pi$. Because the MDP is deterministic, the expectation is taken only over the agent’s policy-induced action sequence. The optimal value of a state $s$ is:
$V(s)\coloneqq\max_{a\in\mathcal{A}}Q(s,a).$
Extensions to Symbolic Regression. In our symbolic regression setting, a state corresponds to a sequence of production rules that define a partially constructed expression, or equivalently, to the class of mathematical expressions that can be generated by completing this partially-complete expression. An action represents the selection of an available production rule that extends the current expression. The state space is the set of all possible expressions of maximum length $H$, and the action space is the set of production rules.
We further provide concrete instantiations of this MDP for our Monte Carlo tree search implementation in Appendix B.4.1 and for deep reinforcement learning in Appendix B.4.2.
B.2 Proof of Theorem 3.1
B.2.1 Problem settings
A planning algorithm is any procedure that, given a model of the MDP and a limited computational budget, uses simulated rollouts to estimate the value of available actions and to recommend a single action for execution. Suppose that after $n$ simulations of a planning algorithm, the agent recommends an action $a_{n}$ to execute at the current state $s$. Its estimated value is $Q(s,a_{n})$, obtained by evaluating the discounted return that results from first taking $a_{n}$ and then following an optimal policy for the remaining steps. The simple regret of this recommendation is
$\mathtt{regret}(n)\coloneqq V(s)-Q(s,a_{n}),$ (6)
which quantifies the loss in discounted return incurred by choosing $a_{n}$ instead of the truly optimal action. A smaller $\mathtt{regret}(n)$ indicates that the planning algorithm used its limited simulation Budget more effectively to identify a near-optimal action.
@@ B.2.1 Problem settings
In standard MCTS, the root of the search tree corresponds to the initial state $s_{0}\in S$. A node at depth $h$ represents an action sequence $a=(a_{1},\ldots,a_{h})\in\mathcal{A}^{h}$, which leads to a terminal state $s(a)$. The search tree at iteration $n$ is denoted $\mathtt{tree}(T_{n})$, and its maximum expanded depth is $d_{n}$.
Historical research has shown that the UCT principle suffers from a theoretical worst case in which it is difficult to derive meaningful regret bounds, making it unsuitable for analyzing the benefit of Egg over standard MCTS. Therefore, we assume that MCTS operates under the Optimistic Planning for Deterministic Systems (OPD) framework [DBLP:conf/acml/LeurentM20] rather than the UCT principle.
B.2.2 Regret bound of MCTS
@@
$\mathtt{regret}(n)=\widetilde{\mathcal{O}}\!\left(n^{-\frac{\log(1/\gamma)}{\log\kappa}}\right).$ (8)
Here, the soft-O version of big-O notation for time complexity is defined as: $f_{n}=\widetilde{\mathcal{O}}(n^{-\alpha})$ means that decays at least as fast as $n^{-\alpha}$, up to a polylog factor.
@@
The theoretical analysis of Egg-MCTS is built on top of the analysis procedure in [DBLP:conf/acml/LeurentM20]. Unlike their analysis, which requires unrolling the graph into a tree and addressing potentially infinite-length paths induced by cycles, our analysis starts from the unrolled tree derived from the graph in [DBLP:conf/acml/LeurentM20].
@@
Let $\mathtt{tree}(T)$ be the search tree after $n$ expansions. Each node corresponds to an action sequence $(a_{0},a_{1},\dots,a_{H-1})$ from the root, and let $s$ denote the MDP state reached by executing this sequence from the root state $s_{0}$. The true value associated with this node is the optimal value of the reached state, denoted by $V(s)$. A pair of functions $(L_{n},U_{n})$ is said to provide lower and upper bounds for $V$ on $\mathtt{tree}(T)$ if
$L_{n}\leq V(s)\leq U_{n},\qquad\text{ for internal node $s$ in the }\mathtt{tree}(T)$
@@
We define the near-optimal branching factor associated with bounds $(L_{n},U_{n})$ as
$\kappa(L_{n},U_{n})\coloneqq\limsup_{h\to\infty}\big|T_{h}^{\infty}(L_{n},U_{n})\big|^{1/h}\in[1,K],$ (9)
@@
Since $(L_{n},U_{n})$ tighten with $n$, the sequence of bounds $\bigl(L_{n},U_{n}\bigr)_{n\geq 0}$ is non-increasing, i.e.,
$0\leq\cdots\leq L_{n-1}\leq L_{n}\leq V\leq U_{n}\leq U_{n-1}\leq\cdots\leq V_{\max}$
they will finally converges to a limit $\kappa_{\infty}=\lim_{n\to\infty}\kappa(L_{n},U_{n})$.
B.2.3 Regret bound of Egg-MCTS
@@
Let $\kappa_{n}=\kappa(L_{n},U_{n})$, and define $\kappa_{\infty}=\lim_{n\to\infty}\kappa(L_{n},U_{n})\in[1,K]$. Then Egg-MCTS achieves the regret bound
@@
The analysis is similar to that of DBLP:conf/acml/LeurentM20. In their approach, the graph is unrolled into a tree to leverage the analysis of Theorem B.1. This unrolled tree contains every action sequence that can be traversed in $G_{n}$. There would exist a path of infinite length if the graph contains cycles. The behavior of the graph $G_{n}$ is therefore analyzed through its corresponding unrolled tree $\mathtt{UnrollTree}(T_{n})$.
Our Egg-MCTS behaves almost identically to this unrolled tree, except that it contains no paths of infinite length, since our algorithm cannot empirically generate or update along an infinite action sequence. Transporting the graph-based value bounds $(L_{n},U_{n})$ to the unrolled tree and applying the depth–regret argument of DBLP:conf/acml/LeurentM20 yields the rate in (10). The detailed derivation is therefore omitted here for brevity. ∎
@@
B.3 Proof of Theorem 3.2
B.3.1 Notation and Assumptions
@@
B.3.2 Proof of Unbiasedness
@@
$\mathbb{E}\;\left[\,\nabla_{\theta}\log p_{\theta}(\tau)\,|\,\phi\right]=\sum_{\tau\in\mathcal{S}_{\phi}}\nabla_{\theta}\log p_{\theta}(\tau)\;\frac{p_{\theta}(\tau)}{q_{\theta}(\phi)}=\frac{1}{q_{\theta}(\phi)}\sum_{\tau\in\mathcal{S}_{\phi}}\nabla_{\theta}p_{\theta}(\tau)$ $=\frac{1}{q_{\theta}(\phi)}\nabla_{\theta}\sum_{\tau\in\mathcal{S}_{\phi}}p_{\theta}(\tau)$
@@
where we used class-invariance of the reward and (11). This equals the expression obtained for $g_{\mathtt{egg}}$, proving unbiasedness. ∎
B.3.3 Proof of Variance Reduction
@@
$\mathbb{E}\;\left[\,Z|\phi\right]=\mathtt{reward}(\phi)\,\mathbb{E}\;\left[\,\nabla_{\theta}\log p_{\theta}(\tau)|\phi\right]=\mathtt{reward}(\phi)\,\nabla_{\theta}\log q_{\theta}(\phi)=g_{\mathtt{egg}}(\theta).$
Thus, $g_{\mathtt{egg}}(\theta)$ is the Rao–Blackwellization of $g(\theta)$ with respect to $\sigma(\phi)$ [8a19a051-2be1-32f8-8b83-075ff1f85473].
@@
$\mathbb{V}\mathrm{ar}_{\tau\sim p_{\theta}}\big[g_{\mathtt{egg}}(\theta)\big]\leq\mathbb{V}\mathrm{ar}_{\tau\sim p_{\theta}}\big[g(\theta)\big]$.
@@
$\mathbb{V}\mathrm{ar}(Z)=\mathbb{V}\mathrm{ar}\big(\mathbb{E}[Z|\phi]\big)\;+\;\mathbb{E}\big[\mathbb{V}\mathrm{ar}(Z|\phi)\big]\;\geq\;\mathbb{V}\mathrm{ar}\;\big(\mathbb{E}[Z|\phi]\big).$
Substituting $Z=g(\theta)$ and $\mathbb{E}[Z|\phi]=g_{\mathtt{egg}}(\theta)$ yields
$\mathbb{V}\mathrm{ar}_{\tau\sim p_{\theta}}\big[g_{\mathtt{egg}}(\theta)\big]\leq\mathbb{V}\mathrm{ar}_{\tau\sim p_{\theta}}\big[g(\theta)\big]$
@@
B.4 Implementation of Egg-embedded Symbolic Regression
B.4.1 Implementation of Egg-MCTS
Markov Decision Process Definition for MCTS. We model the search process as a finite-horizon Markov decision process (MDP). A state corresponds to a partially constructed expression, and each node in the MCTS search tree represents one such state. An action corresponds to the application of a single production rule to the first non-terminal symbol in the current state. The transition dynamics are deterministic: given a state–action pair, the successor state is uniquely determined by the application of the selected production rule. Rewards are assigned only at terminal states (leaf nodes of the search tree), obtained by evaluating the completed expression on the dataset according to an evaluation metric. The episode length is thus equal to the number of applied production rules, and we set the discount factor to $1$. A trajectory is a sequence of state–action transitions corresponding to a path from the root node to a leaf node in the search tree, resulting in either a valid or an invalid expression. The MDP horizon is finite: each episode terminates when the path from the root corresponds to a valid expression with no non-terminal symbols or when the maximum depth $H$ is reached.
Example B.1.
Figure 7 illustrates an example execution pipeline of Egg-MCTS. We focus on two distinct root-to-leaf paths in the search tree:
$\hskip-10.00002pt\text{Path 1: }(A\to A+A,\,A\to\sin^{2}(x_{2}),\,A\to A+A,\,A\to\cos^{2}(x_{2})),\hfill\\ \hfill\text{Node $s_{1}$: }\sin^{2}(x_{2})+\cos^{2}(x_{2})+A.\\ \text{Path 2: }(A\to A+A,\,A\to A/A,\,A\to x_{1},\,A\to x_{1}),\hfill\\ \hfill\text{Node $s_{2}$: }x_{1}/x_{1}+A.\hskip 50.00008pt$
Under the grammar in Section 2, $s_{1}$ corresponds to the partially specified expression $\sin^{2}(x_{2})+\cos^{2}(x_{2})+A$, whereas $s_{2}$ corresponds to $x_{1}/x_{1}+A$. The two nodes are symbolically equivalent under the rewrite rules $\{\sin^{2}(a)+\cos^{2}(a)\leadsto 1,\;a/a\leadsto 1\}$, and both reduce to the same partially specified form $1+A$.
Figure 7: Two distinct paths reach leaf nodes that represent the same underlying function $\phi=1+A$, where AA denotes an arbitrary sub-expression. This equivalence follows from the rewrite rules $\sin^{2}(a)+\cos^{2}(a)\leadsto 1$ and $a/a\leadsto 1$.
Consequently, their rewards—which estimate the average goodness-of-fit on the dataset—should be approximately identical. Exploring the subtrees rooted at $s_{1}$ and $s_{2}$ independently, therefore, incurs redundant computation. In contrast, Egg identifies their equivalence and jointly updates the visit counts and reward estimates along the two paths (blue and orange), eliminating the need to further explore duplicate subtrees in subsequent iterations.